# Retrieve, Summarize, Plan: Advancing Multi-hop Question Answering with an Iterative Approach

### v1

@@ Abstract
Retrieve, Summarize, Plan: Advancing Multi-hop Question Answering with an Iterative Approach
Zhouyu Jiang , Mengshu Sun , Lei Liang , Zhiqiang Zhang Ant Group {jiangzhouyu.jzy,mengshu.sms,leywar.liang,lingyao.zzq}@antgroup.com
1 Introduction
@@
With the rapid development of large language models (LLMs), retrieval-augmented generation (RAG) based on these LLMs has become a popular method for addressing open-domain question answering (Siriwardhana et al., 2023; Lewis et al., 2020; Shi et al., 2024). The typical RAG process involves using a retriever to recall documents from a corpus that are relevant to a given query and using these documents as context inputs for the LLMs to generate responses. However, when dealing with multi-hop question answering, conventional RAG techniques frequently fall short of aggregating all the critical information within a singular retrieval iteration, leading to incomplete or incorrect answers. Consequently, a new genre of iterative RAG methods that incorporate question planning has recently been developed (Shao et al., 2023; Trivedi et al., 2023; Asai et al., 2024). These methods assess after each retrieval whether the information at hand is sufficient for answering the question. If it is not, the methods generate a sub-question for the next step and perform another retrieval, iterating this process until the question can be satisfactorily answered. Owing to the employment of multiple retrieval iterations, iterative RAG has achieved a notable improvement in multi-hop question-answering scenarios compared to the single-round RAG approaches.
@@
Our experimental findings reveal that, under uniform experimental settings, the ReSP model markedly surpasses a range of current single-round and iterative RAG approaches when evaluated on multi-hop question-answering benchmarks such as HotpotQA (Yang et al., 2018) and 2WikiMultihopQA (Ho et al., 2020). Notably, it exhibits a substantial enhancement in performance, with an increase of 4.1 F1 score over the state-of-the-art (SOTA) on HotpotQA, and an improvement of 5.9 F1 score over the SOTA on 2WikiMultihopQA. Furthermore, we conducted a series of in-depth comparative studies to discern the effect of model size on its performance and confirmed that ReSP possesses commendable robustness to context length compared to other RAG methods.
@@ 2 Related Works
Retrieval-Augmented Generation. Retrieval-Augmented Generation (RAG) enhances LLMs by retrieving relevant documents from external databases and integrating them into the generation process (Lewis et al., 2020; Khandelwal et al., 2020; Izacard and Grave, 2021; Jiang et al., 2024). Recent work can be divided into single-round RAG (Kim et al., 2024; Xu et al., 2024; Shi et al., 2024) and iterative RAG (Trivedi et al., 2023; Shao et al., 2023; Jiang et al., 2023; Asai et al., 2024) based on the number of retrieval rounds. In multi-hop question-answering scenarios, iterative RAG often achieves better results because it allows for detailed decomposition of the question. However, due to the increased number of iterations, iterative RAG faces challenges in long-context processing.
@@
Figure 1: The ReSP framework consists of four modules: Reasoner, Retriever, Summarizer, and Generator. The reasoner makes decisions based on the current memory queues, determining whether to exit the iteration and generate a response or to produce a sub-question for further iteration. The retriever searches the corpus based on the sub-question provided by the reasoner (for the first iteration, the sub-question is the same as the overarching question, thus the reasonser is bypassed). The summarizer performs dual summarization on the retrieved documents, extracting information relevant to both the overarching question Q and the current sub-question Q*, and stores the summaries in the global evidence memory and local pathway memory queues respectively. The generator produces answer A based on the information in the memory queues.
3 Methodology
@@
Figure 1 illustrates our ReSP framework, which consists of four components: Reasoner, Retriever, Summarizer, and Generator. Essentially, Reasoner, Summarizer, and Generator are all based on a fine-tuning-free LLM, designed to execute specific tasks using an array of carefully selected prompt engineering techniques. For specific prompt templates, please see Appendix A. Our main contribution lies in the Summarizer, while the design of the other modules is similar to that of conventional iterative RAG methods.
3.1 Dual-Function Summarizer
@@
3.2 Summary-Enhanced Iterative RAG Process
Here we delineate the refined iterative RAG workflow that incorporates the dual-function summarizer.
@@ 3.2 Summary-Enhanced Iterative RAG Process
4 Experimental Settings and Results
4.1 Datasets
We conduct experiments on two multi-hop question-answering benchmark datasets: HotpotQA (Yang et al., 2018) and 2WikiMultihopQA (Ho et al., 2020). Following the open-sourced RAG toolkit FlashRAG (Jin et al., 2024), we employ its preprocessed dataset format. For each dataset, we utilize the first 1,000 entries from the original development set for testing. We report the token-level F1 score of answer strings to evaluate the quality of the generation.
4.2 Experimental Setup
In our main experiment, we employ the Llama3-8B-instruct (AI@Meta, 2024) as the base model and the E5-base-v2 (Wang et al., 2022) as the retriever, while utilizing Wikipedia data from December 2018 as the retrieval corpus. For the model and data links, please refer to the FlashRAG open-source repository  [^1] https://github.com/RUC-NLPIR/FlashRAG.
@@ 4.3 Baselines
Iterative RAG: Iter-RetGen (Shao et al., 2023) leverages the model output from the previous iteration as a specific context to help retrieve more relevant knowledge. IRCoT (Trivedi et al., 2023) guides the retrieval with Chain-of-Thought (CoT) (Wei et al., 2022) and in turn uses retrieved results to improve CoT.
4.4 Main Results
Method Pipeline HotpotQA 2Wiki
Standard RAG Single 38.6 20.1
SuRe Single 33.4 20.6
RECOMP Single 37.5 32.4
REPLUG Single 31.2 21.1
Iter-RetGen Iterative 38.3 21.6
IRCoT Iterative 43.1 32.4
ReSP(ours) Iterative 47.2 38.3
Table 1: Performance comparison on HotpotQA and 2WikiMultihopQA. We report the token-level F1 score of answer strings. All methods utilize fine-tuning-free Llama3-8B-instruct for generation.
@@ 4.4 Main Results
Our method, ReSP, achieves significant improvements on both datasets, outperforming the SOTA by 4.1 F1 score on HotpotQA and 5.9 F1 score on 2WikiMultihopQA, surpassing a range of existing iterative RAG methods. This demonstrates the effectiveness of the approach we propose.
Reasoner Summarizer Generator HotpotQA 2Wiki
@@
Table 2: Impact of base model size on different modules.
5 Empirical Analysis
To further analyze ReSP, we conduct comparative experiments to investigate the impact of the base model size on modules’ performance. Additionally, we examine the robustness of the method to context length. Through case studies, we demonstrate that ReSP can mitigate the issues of over-planning and repetitive planning.
5.1 Impact of the base model size
@@
5.2 Robustness to context length
To determine whether ReSP can address the issue of context overload, we adjust the number of documents retrieved per iteration and observe the changes in performance. Our comparison involves using standard RAG and IRCoT as control groups. These experiments are carried out on the HotpotQA dataset. Results are shown in Figure 2.
Figure 2: Bar chart of the performance variations of different RAG methods with varying numbers of retrieved documents per iteration. We report the token-level F1 score obtained from testing on the HotpotQA dataset.
Firstly, when the number of documents retrieved per iteration k is set to 3, all three methods experience a notable decrease in performance. This indicates that the adequacy of information retrieved in a single round significantly affects both single-round and iterative RAG methods. Therefore, it is necessary to extend the context length during application appropriately.
@@ 5.2 Robustness to context length
"The text does not mention Twisted Fortune, it talks about Eddie Murphy’s career and his brother Charlie Murphy. What’s the next step?"
@@
[Global Evidence Memory]: "Victor Varnado directed the black comedy Twisted Fortune, which starred Charlie Murphy, a brother of Eddie Murphy."
–>[Judgement]: "Yes"
–>[Response]: "Charlie Murphy"
@@
"Rachelle Amy Beinart is not mentioned in any of the provided Wikipedia articles."
@@
"Rachelle Amy Beinart is not mentioned."
@@
[Sub-question]: "Who is Rachelle Amy Beinart?"
@@
[Sub-question]: "What film mentioned in the provided passages features a group of rebels on a mission?"
Table 3: Case studies comparing IRCoT and ReSP. They are conducted to provide evidence of ReSP’s capability in addressing the issues of over-planning and repetitive planning.
5.3 Case Study

### v2

@@ Abstract.
Retrieval-Augmented Generation, Question Answering, LLMs
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–08; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Question answering
1. Introduction
@@
With the rapid development of large language models (LLMs), retrieval-augmented generation (RAG) based on these LLMs has become a popular method for addressing open-domain question answering (Siriwardhana et al., 2023; Lewis et al., 2020; Liu et al., 2023; Cai et al., 2024; Qin et al., 2023). The typical RAG process involves using a retriever to recall documents from a corpus that are relevant to a given query and using these documents as context inputs for the LLMs to generate responses. However, when dealing with multi-hop question answering, conventional RAG techniques frequently fall short of aggregating all the critical information within a singular retrieval iteration, leading to incomplete or incorrect answers. Consequently, a new genre of iterative RAG methods that incorporate question planning has recently been developed (Shao et al., 2023; Trivedi et al., 2023; Asai et al., 2024). These methods assess after each retrieval whether the information at hand is sufficient for answering the question. If it is not, the methods generate a sub-question for the next step and perform another retrieval, iterating this process until the question can be satisfactorily answered. Owing to the employment of multiple retrieval iterations, iterative RAG has achieved a notable improvement in multi-hop question-answering scenarios compared to the single-round RAG approaches.
Figure 1. RAG pipelines illustration and challenges faced.
@@
Our experimental findings reveal that, under uniform experimental settings, the ReSP model markedly surpasses a range of current single-round and iterative RAG approaches when evaluated on multi-hop question-answering benchmarks such as HotpotQA (Yang et al., 2018) and 2WikiMultihopQA (Ho et al., 2020). Notably, it exhibits a substantial enhancement in performance, with an increase of 4.1 F1 score over the state-of-the-art (SOTA) on HotpotQA, and an improvement of 4.4 F1 score over the SOTA on 2WikiMultihopQA. Furthermore, we conducted a series of in-depth comparative studies to discern the effect of base model choice on its performance and confirmed that ReSP possesses commendable robustness to context length compared to other RAG methods.
@@ 2. Related Works
Retrieval-Augmented Generation. Retrieval-Augmented Generation (RAG) enhances LLMs by retrieving relevant documents from external databases and integrating them into the generation process (Lewis et al., 2020; Khandelwal et al., 2020; Izacard and Grave, 2021; Jiang et al., 2024; Zhou et al., 2024). Recent work can be divided into single-round RAG (Kim et al., 2024; Xu et al., 2024; Shi et al., 2024) and iterative RAG (Trivedi et al., 2023; Shao et al., 2023; Jiang et al., 2023; Asai et al., 2024) based on the number of retrieval rounds. In multi-hop question-answering scenarios, iterative RAG often achieves better results because it allows for detailed decomposition of the question. However, due to the increased number of iterations, iterative RAG faces challenges in long-context processing.
Query-focused summarization. Query-focused summarization (QFS) aims to extract or generate a summary of an input document that directly answers or is relevant to a given query(Dang, 2006). Due to the higher requirements of QFS tasks regarding annotation accuracy and data diversity compared to typical summarization tasks, data scarcity has long-limited performance improvement. Previous work has mainly focused on dataset construction (Nema et al., 2017; Baumel et al., 2016; Zhong et al., 2021; Xu et al., 2023), proposing specialized datasets for scenarios such as news, debates, and meetings.
Recent work has attempted to introduce QFS into single-round RAG for refined information compression (Xu et al., 2024; Edge et al., 2024)but has not extended to iterative RAG, nor has it decomposed the functions of QFS. To the best of our knowledge, we are the first to integrate QFS with iterative RAG to address multi-hop question answering.
@@
Figure 2. The ReSP framework consists of four modules: Reasoner, Retriever, Summarizer, and Generator. The reasoner makes decisions based on the current memory queues, determining whether to exit the iteration and generate a response or to produce a sub-question for further iteration. The retriever searches the corpus based on the sub-question provided by the reasoner (for the first iteration, the sub-question is the same as the overarching question, thus the reasonser is bypassed). The summarizer performs dual summarization on the retrieved documents, extracting information relevant to both the overarching question Q and the current sub-question Q*, and stores the summaries in the global evidence memory and local pathway memory queues respectively. The generator produces answer A based on the information in the memory queues.
3. Preliminary
In this section, we formulate the task of single-round RAG and iterative RAG.
3.1. Single-round RAG
Given a question $q$ and a retrieval corpus $D=\{{d_{i}\}}_{i=1}^{|D|}$, the retrieval-augmented LLM needs to generate an answer $y$ based on the retrieved documents relevant to the question. The process can be represented as:
(1) $y={\rm LLM_{Gen}}([q,D_{q}],\rm Prompt_{Gen})$
where $D_{q}$ is the retrieved documents for the question $q$. $\rm Prompt_{Gen}$ is the designated prompt for QA task. $\rm LLM_{Gen}$ is the role of LLM, which focuses on QA task.
3.2. Iterative RAG
Compared to single-round RAG, iterative RAG introduces an evaluation and decision-making process for the current state in each iteration. Specifically, in each iteration, the LLM will provide an action $a$, which has only binary values $\rm(exit,continue)$, and can be represented as:
(2) $a={\rm LLM_{judge}}([q,{state}],\rm Prompt_{judge})$
where $state$ refers to the historical retrieved documents or the information derived from them. $\rm Prompt_{judge}$ is the designated prompt for judgment. $\rm LLM_{judge}$ is an LLM used to assess the state and provide action signals.
The next step is determined based on the value of $a$:
(3) $next\_step=\begin{cases}{\rm LLM_{plan}}([q,{state}],{\rm Prompt_{plan}}),a={\rm continue}\\ {\rm answer\_generation},a={\rm exit}\end{cases}$
where $\rm Prompt_{plan}$ is the designated prompt for sub-question generation. $\rm LLM_{plan}$ is an LLM for breaking down problems.
4. Methodology
@@
Figure 2 illustrates our ReSP framework, which consists of four components: Reasoner, Retriever, Summarizer, and Generator. Essentially, Reasoner, Summarizer, and Generator are all based on a fine-tuning-free LLM, designed to execute specific tasks using an array of carefully selected prompt engineering techniques. For specific prompt templates, please see Table 1. Our main contribution lies in the Summarizer, while the design of the other modules is similar to that of conventional iterative RAG methods.
Table 1. The prompt templates of different modules in ReSP. The input parameters include Overarching question: the input question; Sub-question: the current sub-question generated by the reasoner; Combined memory queues: the concatenated content of the global evidence memory and local pathway memory queues for each iteration; docs: the documents retrieved from a single round of retrieval.
Module Function Prompt
[…]
4.1. Dual-Function Summarizer
@@
4.2. Summary-Enhanced Iterative RAG Process
Here we delineate the refined iterative RAG workflow that incorporates the dual-function summarizer. The pseudocode for the algorithm implementation is shown as Algorithm 1.
@@ 4.2. Summary-Enhanced Iterative RAG Process
Algorithm 1 Multi-hop QA through ReSP
0:  Overarching question $Q$, Document corpus $D$, Retriever $R$, LLMs with different roles $(LLM_{Reason},LLM_{Sum},LLM_{Gen})$, Max number of iterations $M$
  $if\_finished\leftarrow{\rm False}$
  $run\_cnt\leftarrow 0$
  $q\leftarrow Q$
  $S_{Global}\leftarrow[]$
  $S_{Local}\leftarrow[]$
  while not $if\_finished$ and $run\_cnt<M$ do
     $run\_cnt\leftarrow run\_cnt+1$
     $D_{q}\leftarrow R(q)$
     if $q\neq Q$ then
        $S_{Local}\leftarrow S_{Local}+LLM_{Sum}([q,D_{q}],{\rm Prompt_{local}})$
     end if
     $S_{Global}\leftarrow S_{Global}+LLM_{sum}([Q,D_{q}],{\rm Prompt_{global}})$
     $if\_finished\leftarrow LLM_{Reason}([Q,S_{Global},S_{Local}],{\rm Prompt_{judge}})$
     if not $if\_finished$ then
        $Q^{*}\leftarrow LLM_{Reason}([Q,S_{Global},S_{Local}],{\rm Prompt_{plan}})$
        $q\leftarrow Q^{*}$
     end if
  end while
  $A\leftarrow LLM_{Gen}([Q,S_{Global},S_{Local}],{\rm Prompt_{Gen}})$
  Final answer $A$
5. Experimental Settings and Results
5.1. Datasets
We conduct experiments on two multi-hop question-answering benchmark datasets: HotpotQA (Yang et al., 2018) and 2WikiMultihopQA (Ho et al., 2020). These two datasets are both constructed based on Wikipedia documents, which enables the utilization of a consistent document corpus and retrievers to provide external references for LLMs. Following the open-sourced RAG toolkit FlashRAG (Jin et al., 2024), we employ its preprocessed dataset format. For each dataset, we utilize the first 1,000 entries from the original development set for testing. We report the token-level F1 score of answer strings to evaluate the quality of the generation.
5.2. Experimental Setup
In our main experiments, we employ the Llama3-8B-instruct(AI@Meta, 2024) as the base model and E5-base-v2 (Wang et al., 2022) as the retriever. The retrieval corpus comprises Wikipedia data from December 2018, consisting of approximately 20 million pre-segmented passages. For the model and data links, please refer to the FlashRAG open-source repository  [^1] https://github.com/RUC-NLPIR/FlashRAG.
@@ 5.3. Baselines
Direct Prompting: LLM directly generates answers without retrieval.
@@
Iterative RAG: Iter-RetGen (Shao et al., 2023) leverages the model output from the previous iteration as a specific context to help retrieve more relevant knowledge. Self-RAG (Asai et al., 2024) enhances an LM’s quality and factuality through retrieval and self-reflection. FLARE (Jiang et al., 2023) iteratively uses the upcoming sentence to anticipate future content, which is then utilized as a query to retrieve relevant documents to regenerate the sentence if it contains low-confidence tokens. IRCoT (Trivedi et al., 2023) guides the retrieval with Chain-of-Thought (CoT) (Wei et al., 2022) and in turn uses retrieved results to improve CoT.
5.4. Main Results
Table 2. Performance comparison on HotpotQA and 2WikiMultihopQA. All methods except Self-RAG utilize fine-tuning-free Llama3-8B-instruct for generation. Self-RAG uses trained selfrag-llama2-7B released by authores. We run experiments five times and report the average token-level F1 score of answer strings. The result of ReSP outperforms baseline models in t-test at p ¡ 0.05 level. The best results are in bold and the second best results are underlined.
Method Pipeline type HotpotQA 2WikiMultihopQA
Direct Prompting 28.4 33.9
Standard RAG Single-round RAG 38.6 20.1
SuRe (Kim et al., 2024) Single-round RAG 33.4 20.6
RECOMP(Xu et al., 2024) Single-round RAG 37.5 32.4
REPLUG(Shi et al., 2024) Single-round RAG 31.2 21.1
Iter-RetGen(Shao et al., 2023) Iterative RAG 38.3 21.6
Self-RAG(Asai et al., 2024) Iterative RAG 29.6 25.1
FLARE (Jiang et al., 2023) Iterative RAG 28.0 33.9
IRCoT(Trivedi et al., 2023) Iterative RAG 43.1 32.4
ReSP(ours) Iterative RAG 47.2 38.3
@@ 5.4. Main Results
Our method, ReSP, achieves significant improvements on both datasets, outperforming the SOTA by 4.1 F1 score on HotpotQA and 4.4 F1 score on 2WikiMultihopQA, surpassing a range of existing iterative RAG methods. This demonstrates the effectiveness of the approach we propose.
To analyze whether ReSP can address context overload and redundant planning issues, we record key statistics during the model’s execution.
Initially, concerning the issue of context overload, utilizing HotpotQA as an example, when the maximum number of iterations is limited to three, the recorded average total word count of the retrieved documents during the process is 509.59, while the average word count of the evidence information provided to the generator is 86.61, culminating in a compression ratio of 5.88. This indicates that ReSP significantly compresses useful information during the process, helping to prevent context overload.
Regarding redundant planning, ReSP exhibits an average iteration count of 1.24 for HotpotQA and 1.60 for 2WikiMultihopQA, while IRCoT has counts of 1.80 and 2.32, respectively. This indicates that ReSP considerably minimizes planning steps.
Additionally, as shown in Table 3, we provide a comparison of computation time. We record the runtime of standard RAG, IRCoT, and ReSP on 200 instances, maintaining the same experimental environment described before. The results indicate that the iterative ReSP takes longer compared to the single-round standard RAG, which may make it less suitable for time-sensitive scenarios unless specialized engineering optimizations are performed. However, compared to IRCoT, which is also an iterative method, ReSP significantly reduces runtime by compressing context and minimizing redundant planning while achieving better performance.
It is noteworthy that ReSP allows for the exposure of both global and local evidence information collected at each step of the problem-solving process. This is particularly suitable for scenarios that demand high transparency and trustworthiness of model outputs, such as legal and medical fields. In these contexts, users can understand how the problem is decomposed, how information is retrieved, and how the final answer is organized during the process. This advantage is not present in standard RAG and other iterative RAG methods that do not record the retrieval process.
Table 3. Comparison of computation time.
Method HotpotQA 2WikiMultihopQA
Standard RAG 1.1min 1.2min
IRCoT 41.8min 48.2min
ReSP 9.5min 13.7min
Table 4. Impact of base model size on different modules.
Reasoner Summarizer Generator HotpotQA 2WikiMultihopQA
@@
6. Empirical Analysis
To further analyze ReSP, we conduct comparative experiments to investigate the adaptability across different open-source base models and the impact of the base model size on modules’ performance. Additionally, we examine the method’s robustness concerning two key variables: context length and maximum number of iterations. Through case studies, we demonstrate that ReSP can mitigate the issues of over-planning and repetitive planning.
6.1. Adaptability across different base models
Considering popularity, performance, and computational cost, we utilized the open-source Llama3-8B-Instruct as the base model in our main experiments to validate the effectiveness of our method. In practical scenarios, there are many other base model series to choose from. To assess the adaptability of our method across different base model series, we select three open-source base models of similar size to Llama3-8B-Instruct for experiments, maintaining consistency in other experimental settings: Qwen2.5-7B-Instruct(Team, 2024), gemma-2-9b-it(Jeanine Banks, 2024), and Deepseek-V2-Lite-Chat(DeepSeek-AI, 2024). All models can be available on HuggingFace [^2] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct [^3] https://huggingface.co/google/gemma-2-9b-it [^4] https://huggingface.co/deepseek-ai/DeepSeek-V2-Lite-Chat.
The results on HotpotQA are shown in Figure 3. It can be observed that, compared to Standard RAG and IRCoT, the ReSP method consistently achieves significant improvements across three different base models. This demonstrates our method’s strong adaptability to various base models.
Figure 3. Performance comparison across different base models. We report the token-level F1 score obtained from testing on the HotpotQA dataset.
6.2. Impact of the base model size
@@
6.3. Robustness to context length
Figure 4. Bar chart of the performance variations of different RAG methods with varying numbers of retrieved documents per iteration. We report the token-level F1 score obtained from testing on the HotpotQA dataset.
To further determine whether ReSP can address the issue of context overload, we adjust the number of documents retrieved per iteration and observe the changes in performance. Our comparison involves using standard RAG and IRCoT as control groups. These experiments are carried out on the HotpotQA dataset. Results are shown in Figure 4. Firstly, when the number of documents retrieved per iteration k is set to 3, all three methods experience a notable decrease in performance. This indicates that the adequacy of information retrieved in a single round significantly affects both single-round and iterative RAG methods. Therefore, it is necessary to extend the context length during application appropriately.
Table 5. Case studies comparing IRCoT and ReSP. They are conducted to provide evidence of ReSP’s capability in addressing the issues of over-planning and repetitive planning.
@@ 6.3. Robustness to context length
”The text does not mention Twisted Fortune, it talks about Eddie Murphy’s career and his brother Charlie Murphy. What’s the next step?”
@@
[Global Evidence Memory]: ”Victor Varnado directed the black comedy Twisted Fortune, which starred Charlie Murphy, a brother of Eddie Murphy.”
–>[Judgment]: ”Yes”
–>[Response]: ”Charlie Murphy”
@@
”Rachelle Amy Beinart is not mentioned in any of the provided Wikipedia articles.”
@@
”Rachelle Amy Beinart is not mentioned.”
@@
[Sub-question]: ”Who is Rachelle Amy Beinart?”
@@
[Sub-question]: ”What film mentioned in the provided passages features a group of rebels on a mission?”
6.4. Exploration of the number of iterations
Figure 5. Performance with different maximum number of iterations.
In our main experiments, the maximum number of allowed iterations was set to three, following the construction methods detailed in the HotpotQA and 2WikiMultihopQA papers, where most multi-hop questions are resolved within three hops. However, in practical applications, the maximum number of iterations can significantly affect the performance of iterative RAG methods. Consequently, we conducted comparative experiments by varying the maximum iterations from 1 to 5 to observe the resulting performance changes in ReSP.
As illustrated in Figure 5, a significant decline in ReSP’s performance is observed when the maximum number of iterations is less than three, indicating that gathering information over multiple rounds can enhance response accuracy. However, when the maximum number of iterations exceeds three, the performance improvements become negligible. This is attributed to ReSP’s optimization for redundant planning, enabling problem resolution within a reasonable number of hops. These findings are consistent with the average iteration statistics presented in the Section 5.4.
6.5. Case Study